0.447 mol of a solid was dissolved in 240 mL of water at 20.6 oC. After the solid had fully dissolved, the final temperature of the solution was 27.1 oC. What is the molar heat of solution of the substance?
Answer #1

Specific heat of water = s = 4.18 J/g°C

Mass of water (m) = volume * density = 240 * 1 = 240 g

Heat of solution (H) = m * s * delta T = 240 * 4.18 * (27.1 - 20.6)

H = 6520.8 J

Molar heat = Heat/moles = 6520.8/0.447 = 14587.92 J/mol

= 14.588 KJ/mol .... Answer

Please let me know in case of any concerns.